# Emancipatory Information Retrieval

### v3

@@ 3.3 Provocations
Emancipatory IR must leverage instruments of provocation to challenge our thinking and help us reimagine radically alternative futures. The inception of emancipatory IR must not be in a vacuum, but must be informed by emancipatory ideologies borne of our historical and ongoing struggles, including feminist, queer, decolonial, anti-racist, anti-casteist, anti-ableist, and abolitionist thoughts. Our challenge is not only to imagine radically new emancipatory futures, but also practicing those desired values and build prefigurative counter-structures both in the form of new technological designs and sociopolitical rearrangements within our communities. The task to change technology is inextricably linked to the task of changing ourselves and our own social organizations. The practice of reflexivity becomes important in this context to articulate (and critique if necessary) our own values, positionality, and incentives, as well as our relationships with institutions and social structures that we in turn intend to change. It is in this process that we desperately need to engage scholars outside of computing, learn from generations of activists who have been engaged in emancipatory struggles, and let ourselves be influenced by the artists and writers that dare to imagine radically free futures and humanistic social arrangements.
@@ 4.3 Where do we begin?
    A futuring session where participants are encouraged to radically reimagine information retrieval systems and how they should be embedded in our society and our daily lives.
@@
    Following the futuring exercise, participants should then sketch out the path to our desired futures and identify concrete challenges and initiatives that the emancipatory IR community should take on. Some of these may align with the projects enumerated in Section 3.2 or may identify the need for additional projects that we have not considered here.
@@
The goal of such a workshop would be to build trust, solidarity, and a shared understanding of our aspirations and plans within the seed community. It is also our aim with this workshop to challenge the seemingly unassailable wall of colonial capitalist values that have arrested our present visions of society and technologies. We hope such a workshop can be successful in finding the cracks in the wall through which we can collectively peep out and free ourselves to imagine the pluralism of social arrangements that are possible free from oppression in which we can all flourish.
@@ 5 Conclusion
This piece is a provocation for the IR community to situate itself in the generational emancipatory struggles and stand in resistance to the growing epidemic of authoritarianism and fascism that we are experiencing worldwide today. While focused on IR, it is a challenge to the broader computing and technology community to reimagine our work to not serve the privileged few by further concentrating immense power and wealth, but as a liberating force that connects us and humanizes us universally. Technology after all is a project of world-making. So, what kind of a world are we collectively going to build together?

### v4

@@ 3.3 Provocations
Emancipatory IR must leverage instruments of provocation to challenge our thinking and help us reimagine radically alternative futures. The inception of emancipatory IR must not be in a vacuum, but must be informed by emancipatory ideologies borne of our historical and ongoing struggles, including feminist, queer, decolonial, anti-racist, anti-casteist, anti-ableist, and abolitionist thoughts. Our challenge is not only to imagine radically new emancipatory futures, but also practicing those desired values and build prefigurative counter-structures both in the form of new technological designs and sociopolitical rearrangements within our communities. The task to change technology is inextricably linked to the task of changing ourselves and our own social arrangements. The practice of reflexivity becomes important in this context to articulate (and critique if necessary) our own values, positionality, and incentives, as well as our relationships with institutions and social structures that we in turn intend to change. It is in this process that we desperately need to engage scholars outside of computing, learn from generations of activists who have been engaged in emancipatory struggles, and let ourselves be influenced by the artists and writers that dare to imagine radically free futures and humanistic social structures.
@@ 4.3 Where do we begin?
    A design futuring session where participants are encouraged to radically reimagine IR systems and how they should be embedded in our society and our daily lives.
@@
    Following the futuring exercise, participants should then sketch out paths to our desired futures and identify concrete challenges and initiatives that the emancipatory IR community should take on. Some of these may align with the projects enumerated in Section 3.2 or may identify the need for additional projects that we have not considered here.
@@
The goal of such a workshop would be to build trust, solidarity, and a shared understanding of our aspirations and plans within the seed community. It is also our aim with this workshop to challenge the seemingly unassailable wall of colonial capitalist values that have arrested our present visions of society and technologies. We hope such a workshop can be successful in finding the cracks in the wall through which we can collectively peep out and free ourselves to imagine a plurality of social arrangements that are possible in which we can all flourish free from oppression.
@@ 5 Conclusion
The field of IR has undergone seismic shifts over the last decade largely instigated by rapid developments in deep learning [94] and more recently generative AI [137]. Between trying to stay updated in the face of daily firehoses of new AI publications and chasing the coveted badges of “state-of-the-art” it can be easy as an IR researcher to overlook the pressing needs of society and envision only the possibilities as conjured by the latest AI news cycle. If we are to address the multitude of crises facing our world today—spanning the ecological, the social, the political, and the economic spectrums—IR must undergo another seismic shift over the next decade. This piece is a provocation for the IR community to situate itself in the generational emancipatory struggles and stand in resistance to the growing epidemic of authoritarianism and fascism through the practices of making and unmaking of information access technologies. It is also an invitation to scholars from other fields as well as to activists and artists to join us and be part of this movement. While focused on IR, it is also a challenge to the broader computing and technology community to reimagine our work to not serve the privileged few by further concentrating immense power and wealth, but as a liberating force that connects us and humanizes us universally. Technology after all is a project of world-making. So, what kind of a world are we collectively going to build together?